OpenStudy (anonymous):

How do you multiply regular numbers by numbers with radicals. For example, what would 7 times radical 3?

OpenStudy (anonymous):

\[7\times \sqrt{3}=7\sqrt{3}=\sqrt{49*3}=\sqrt{147}\] The last two steps are optional. You usually stop at the second step.

OpenStudy (anonymous):

Thank you so much! I had only gotten to the second part till now.

OpenStudy (anonymous):

Emileokada is correct but you should probably stop at \[7\sqrt{3}\] It is just like multiplying a constant and a variable 3 times y is 3y

OpenStudy (anonymous):

Yeah I usually do, but my work called for me continuing the process and I actually got my problem right because of this.

OpenStudy (anonymous):

ok, follow the directions given. The last part is usually not common since that can be simplified :)